You want to determine the control lines for a "P" chart for quality control purposes. If the total number of defects from all samples is 2,500, the number of samples is 100, and the sample size is 50, which of the following would be the standard deviation used in developing the control lines?

a. .0707

b. .4900

c. .2499

d. .1556

e. .02499

Please show all steps and where all numbers come from if they aren't already given.
Answer #1

1. A. 0707

P-BAR = TOTAL NUMBER OF DEFECTS / (SAMPLE SIZE * NUMBER OF SAMPLES)

STDEV = SQRT((PBAR * (1 - PBAR)) / SAMPLE SIZE

TOTAL DEFECTS = 2500
SAMPLE SIZE = 50
NUMBER OF SAMPLES = 100


PBAR = 2500 / (50 * 100) = 0.5
STDEV = SQRT((0.5 (1 - 0.5)) / 50 = 0.0707